A company currently pays a dividend of $3.25 per share (D0 = $3.25). It is estimated that the company's dividend will grow at a rate of 18% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company's stock has a beta of 1.25, the risk-free rate is 8%, and the market risk premium is 6%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=8+(1.25*6)=15.5%

D1=(3.25*1.18)=3.835

D2=(3.835*1.18)=4.5253

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(4.5253*1.05)/(0.155-0.05)

=45.253

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.835/1.155+4.5253/1.155^2+45.253/1.155^2

=$40.63(Approx).